determine whether these functions are bijections f:Q to R x2+1/x


1
Expert's answer
2021-09-27T12:25:46-0400

Let us determine whether the function "f:\mathbb Q \to \R,\ f(x)= x^2+\frac{1}x," is a bijection.

Let "x\in\mathbb Q," that is "x" be a rational number. Taking into account that the sum, the product and the fraction of two ratianal number is a rational number, we conclude that "x^2\in\mathbb Q,\ \frac{1}x\in\mathbb Q," and hence "f(x)=x^2+\frac{1}x\in\mathbb Q." Therefore, for any "y\in\R\setminus\mathbb Q\subset\R" we get that the preimage "f^{-1}(y)" is emptyset, and thus "f" is not a surjection. Consequently, "f" is not a bijection.